APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday October 24, 2010 (MJD 55494) ---=== OBSERVING TEAM ===--- Steph Snedden (afternoon) Elena Malanushenko Viktor Malanushenko Paul Harding online ---=== OBSERVING PLAN ===--- Engineering and Marvels Science! ---=== OBSERVING SUMMARY ===--- Partly cloudy all the night. At the evening we made Kyle's tests then we run MARVELS observation: We tried to make observations using astrometrical plates. This test could not be completed as there was clouds and up to 2" seeing. We tested the procedure and tested updated codes of guider. Paul was online and cooperated with us. Initial Kyle's procedure needs more then 2 hours. Kyle revised the procedure to make it shorter. We completed this offset test with two plates (ct 11 and ct 16). We did not try focus test because bad seeing. In general impression, this test might be better organized and needs to have more attention to detailes. It could be good to have previous consultations with observers, and prepared some scripts for optimal procedure with developers. We observe 6 plates: Cartridges 6#3706B, 8#3896A, 7#3638A, 9#3635AB, 5#3059A. We aborted scientific exposure of Cart5 plate3059, i.e. it was close to zenith position. So we can not observe that plate close to HA=0. brocken/missed fibers: 6#3706B fibers: 87,88 8#3896A fibers: 47,48 7#3638A fibers: 33,34 9#3635B fibers: 45,46 Tonight we used "guider: trunk+svn116784M". It worked properly, but after disconnection at the end of night we got messages: 12:41:48 .stui 0 stui Disconnected 12:41:48 APO.elenam 259 stui F Aborted; Actor='guider'; Cmd='off'; Text="disconnected" 12:41:48 APO.elenam 265 stui F Aborted; Actor='guider'; Cmd='off'; Text="disconnected" .... it was 29 lines. ---=== OBSERVING LOG ===--- Afternoon: ---------- 19:15Z - BOSS bias, dark, flat and arc were taken in the afternoon after Bob re-installed the HgCd lamps. They looked normal. 22:15Z - Hartmanns taken on cartridge 13, and focus should be good for tonight. No ring adjustments made. All actors responded with normal status, except the guider, and that is because Paul was working on it for tonight's astrometric plate tests. The error at that particular time related to the guiding on decentered stars work. 22:30Z guider 0 guider w text="Unexpected exception type object 'Msg' has no attribute 'DECENTER' in guider master thread" 22:41Z - Elena reported a problem with a minidru monitor last night. It was powered off when I came in. I powered it up and it was fine - no problems. I don't know if anyone worked on it today or not. 23:00Z - Ventilation started. doors opened. Cartridge 11 mounted. Night: ------ Astrometric plate observations from Kyle: Cart 11#3526. Sky was very cloudy, so we could not do good observations. We tried to do this test anyway and to check guider updates with Paul. We used Kyle's email as initial instruction for the test (Kyle, Oct.24, 2010, 3:38:56 PM MDT). Found stars on GTF. #1 Took hartmann and calibration during gotoField. hartmann ff=121140,*41 arc=*42,flat=*43 #2 Turned on guider and centered stars. Took 4 science exposures with 120 sec integration time boss exposure science itime=120 ff=121144-*47 #2 Stopped guider and do another set of hartman and calibrations hartmann ff=121148,*49 arc=*50,flat=*51 #4 Took another four boss exposure boss exposures with integrations 240 sec. ff=121152,53,54,55 #5 stopped gider and took one more set of hartmann and calibration hartmann ff=56,57, flat=58, arc=59 #6 "guider on" guider and centered stars. "guider off" and continue tracking #7 Make offset on Ra on 0.5 arcsec. "tcc offsets arc 0.000139, 0.0" (we work with degrees) Took one boss science exposure itime=120, ff=60 returned to center using opposite Ra offset=-0.000139. Turned "guider on" and centered stars. #8. Did the same as in #7 but with offset Ra -0.5". ff=61 #9 the same as #7 but with Dec offset +0.5", ff=121162 - bad ff=121163 #10 Made offset on Dec -0.5", ff=121164 #11 Return to center, turned guider on, and took one boss exposure with integration 120 sec. It was still a lot of clouds, seeing up to 2.0-2.3", so we postponed the last part of the check focus moving the secondary mirrow. We estimated possible time to make this test and it take more then 2 hours to be completed. Kyle and Paul sent us shorter instruction (that instruction in problem section). Cart 16 plate 3521, astrometric plate Found stars on GTF and made hartmann and calibrations on fly. Guider on and stars were centered. #2 science exposures, itime=120, ff=121170,71 #3 "guider off" but continue to track #4 Applied Ra +0.5" offsets and made sciencse exposure, ff=72 #5 offset telescope back to center and centered stars with guider on. #7,#8 applied the same as #4 and #5 with negative Ra offset -0.5". ff=73 #9,10 - applied the same with Dec=0.5" offset, ff=74 #11,12 - the same with Dec=-0.5" sec offset, ff=75 #12 - took two more science exposures with guider and stars were centered. We did not do focus test because bad seeing and clouds. We switched to Marvels. We did not have right cart availble for us and need to wait ~40 min before to start to work with the next cart as it was at too low Alt. MARVELS Cart 6#3706B 05:17Z Calibration T05:17:19.580, T05:19:27.200, T05:24:02.650 05:30Z T05:30:06.420 6 3706B 3000. 48. 1336. STAR HD219828 Cart 8#3896A 06:23Z Calibration T06:23:01.270, T06:25:08.580, T06:29:45.100 06:42Z T06:42:06.120 8 3896A 3000. 10. 580. STAR HD37605 Cart 7#3638A 07:37Z Calibration T07:37:28.890,T07:39:35.210,T07:44:11.560; clouds 08:05Z Calibration T08:05:09.640,T08:07:17.020,T08:11:54.450 08:14Z T08:14:07.050 7 3638A 3000. 26. 906. STAR HD46375 Cart 9#3635A 09:07Z Calibration T09:07:15.720, T09:09:22.060, T09:13:58.560 09:18Z T09:18:09.570 9 3635A 3000. 31. 965. STAR HD43691 Cart 9#3635B 10:11Z Calibration T10:11:07.400, T10:13:14.760, T10:17:51.220 10:20Z T10:20:11.220 9 3635B 3000. 28. 916. STAR HD49674 Cart 5#3089A 11:12Z Calibration T11:12:50.010, T11:14:57.350, T11:19:35.700 We observed it at the HA=0, telescope was very close to zenith. The tcc motors were halted with: "11:47:27 APO.elenam 572 guider w text="TCC motion failed, but axis motions are bypassed in sop"." We aborted exposure, and finished observations. 12:38Z endNight finished ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 1 19:13 b1-00121121 13 4287 bia 0.0 -12519 -14006 -12345 2 19:15 b1-00121122 13 4287 dar 20.0 -12519 -14006 -12345 3 19:17 b1-00121123 13 4287 fla Out 30.9 -12519 -14006 -12345 4 19:18 b1-00121124 13 4287 arc Out 4.9 -12519 -14006 -12345 5 19:20 b1-00121125 13 4287 arc Out 4.9 -12519 -14006 -12345 6 22:08 b1-00121126 13 4287 arc Lef 4.8 -12519 -14006 -12345 HART 7 22:10 b1-00121127 13 4287 arc Rig 4.8 -12519 -14006 -12345 HART 8 22:19 b1-00121128 13 4287 arc Lef 4.8 -5219 -6706 -5044 HART 9 22:20 b1-00121129 13 4287 arc Rig 4.8 -5219 -6706 -5044 HART 10 00:57 b1-00121130 11 3526 arc Lef 4.8 -5219 -6706 -5044 HART 11 00:58 b1-00121131 11 3526 arc Rig 4.8 -5219 -6706 -5044 HART 12 00:59 b1-00121132 11 3526 arc Out 4.8 -6165 -7656 -5992 13 01:01 b1-00121133 11 3526 fla Out 30.8 -6165 -7656 -5992 14 01:04 b1-00121134 11 3526 bia 0.0 -6165 -7656 -5992 15 01:05 b1-00121135 11 3526 dar 20.0 -6165 -7656 -5992 16 01:08 b1-00121136 11 3526 arc Lef 4.8 -6165 -7656 -5992 HART 17 01:09 b1-00121137 11 3526 arc Rig 4.8 -6165 -7656 -5992 HART 18 01:11 b1-00121138 11 3526 arc Out 4.8 -6157 -7645 -5983 19 01:13 b1-00121139 -1 -1 fla Out 30.9 -6157 -7645 -5983 20 01:16 b1-00121140 11 3526 arc Lef 4.8 -6157 -7645 -5983 HART 21 01:17 b1-00121141 11 3526 arc Rig 4.8 -6157 -7645 -5983 HART 22 01:19 b1-00121142 11 3526 arc Out 4.8 -5997 -7484 -5822 23 01:21 b1-00121143 11 3526 fla Out 30.8 -5997 -7484 -5822 24 01:37 b1-00121144 11 3526 sci Out 120.8 -5997 -7484 -5822 25 01:41 b1-00121145 11 3526 sci Out 120.8 -5997 -7484 -5822 26 01:44 b1-00121146 11 3526 sci Out 120.8 -5997 -7484 -5822 27 01:47 b1-00121147 11 3526 sci Out 120.8 -5997 -7484 -5822 28 01:55 b1-00121148 11 3526 arc Lef 4.8 -5997 -7484 -5822 HART 29 01:57 b1-00121149 11 3526 arc Rig 4.8 -5997 -7484 -5822 HART 30 01:59 b1-00121150 11 3526 fla Out 30.8 -5997 -7484 -5822 31 02:03 b1-00121151 11 3526 arc Out 4.8 -5997 -7484 -5822 32 02:07 b1-00121152 11 3526 sci Out 240.9 -5997 -7484 -5822 33 02:12 b1-00121153 11 3526 sci Out 240.9 -5997 -7484 -5822 34 02:18 b1-00121154 11 3526 sci Out 240.8 -5997 -7484 -5822 35 02:23 b1-00121155 11 3526 sci Out 240.8 -5997 -7484 -5822 36 02:32 b1-00121156 11 3526 arc Lef 4.8 -5997 -7484 -5822 HART 37 02:33 b1-00121157 11 3526 arc Rig 4.8 -5997 -7484 -5822 HART 38 02:35 b1-00121158 11 3526 fla Out 30.8 -5997 -7484 -5822 39 02:40 b1-00121159 11 3526 arc Out 4.8 -5997 -7484 -5822 40 02:45 b1-00121160 11 3526 sci Out 120.8 -5997 -7484 -5822 41 02:51 b1-00121161 11 3526 sci Out 120.8 -5997 -7484 -5822 42 03:00 b1-00121162 11 3526 sci Out 120.8 -5997 -7484 -5822 43 03:06 b1-00121163 11 3526 sci Out 120.8 -5997 -7484 -5822 44 03:12 b1-00121164 11 3526 sci Out 120.8 -5997 -7484 -5822 45 03:18 b1-00121165 11 3526 sci Out 120.8 -5997 -7484 -5822 46 03:42 b1-00121166 16 3521 arc Lef 4.8 -5997 -7484 -5822 HART 47 03:43 b1-00121167 16 3521 arc Rig 4.8 -5997 -7484 -5822 HART 48 03:44 b1-00121168 16 3521 arc Out 4.8 -12836 -14324 -12658 49 03:46 b1-00121169 16 3521 fla Out 30.9 -12836 -14324 -12658 50 03:50 b1-00121170 16 3521 sci Out 120.8 -12836 -14324 -12658 51 03:53 b1-00121171 16 3521 sci Out 120.8 -12836 -14324 -12658 52 03:59 b1-00121172 16 3521 sci Out 120.8 -12836 -14324 -12658 53 04:05 b1-00121173 16 3521 sci Out 120.9 -12836 -14324 -12658 54 04:11 b1-00121174 16 3521 sci Out 120.8 -12836 -14324 -12658 55 04:18 b1-00121175 16 3521 sci Out 120.8 -12836 -14324 -12658 56 04:27 b1-00121176 16 3521 sci Out 120.8 -12836 -14324 -12658 57 04:31 b1-00121177 16 3521 sci Out 120.8 -12836 -14324 -12658 58 04:59 b1-00121178 6 3706 arc Out 4.8 -12836 -14324 -12658 59 05:00 b1-00121179 6 3706 fla Out 30.9 -12836 -14324 -12658 Reading FITS from /data/MARVELS/55494/et1/ - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - Image Cart Plate Texp Exptype Object - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - T04:56:33.520 0 11. 43721. 4818. TIO NULL T04:58:40.940 0 160. 329. 747. THAR NULL T05:03:18.340 0 11. 43842. 4826. TIO NULL T05:17:19.580 0 11. 43879. 4828. TIO NULL T05:19:27.200 0 160. 331. 747. THAR NULL T05:24:02.650 0 11. 43717. 4816. TIO NULL T05:30:06.420 6 3706B 3000. 48. 1336. STAR HD219828 T06:23:01.270 0 11. 45717. 5048. TIO NULL T06:25:08.580 0 160. 344. 762. THAR NULL T06:29:45.100 0 11. 45677. 5054. TIO NULL T06:42:06.120 8 3896A 3000. 10. 580. STAR HD37605 T07:37:28.890 0 11. 44241. 4933. TIO NULL T07:39:35.210 0 160. 336. 755. THAR NULL T07:44:11.560 0 11. 44256. 4938. TIO NULL T08:05:09.640 0 11. 44404. 4951. TIO NULL T08:07:17.020 0 160. 339. 755. THAR NULL T08:11:54.450 0 11. 44289. 4938. TIO NULL T08:14:07.050 7 3638A 3000. 26. 906. STAR HD46375 T09:07:15.720 0 11. 44805. 5038. TIO NULL T09:09:22.060 0 160. 343. 764. THAR NULL T09:13:58.560 0 11. 44985. 5048. TIO NULL T09:18:09.570 9 3635A 3000. 31. 965. STAR HD43691 T10:11:07.400 0 11. 44913. 5045. TIO NULL T10:13:14.760 0 160. 343. 765. THAR NULL T10:17:51.220 0 11. 45132. 5057. TIO NULL T10:20:11.220 9 3635B 3000. 28. 916. STAR HD49674 T11:12:50.010 0 11. 43853. 4890. TIO NULL T11:14:57.350 0 160. 336. 755. THAR NULL T11:19:35.700 0 11. 43893. 4902. TIO NULL - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - - ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:22Z 11#3526 -14.48 0.0000 56.46 0.0000 167.79 0.0000 0.999980 ObjArcOff=0.000337,0.000000,4794690139.84000,0.001228,0.000000,4794690139.84000 05:40Z 6#3706B -61.27 0.0000 65.02 0.0000 128.88 0.0000 1.000110 ObjArcOff=0.002257,0.000000,4794702022.96000,0.001776,0.000000,4794702022.96000 06:54Z 8#3896A 74.00 0.0000 33.37 0.0000 234.40 0.0000 1.000370 ObjArcOff=-0.000852,0.000000,4794706555.53000,0.000902,0.000000,4794706555.53000 10:46Z 9#33635B 130.37 0.0000 76.64 0.0000 -57.65 0.0000 1.000300 ObjArcOff=0.001206,0.000000,4794720481.32000,0.000420,0.000000,4794720481.32000 11:35Z 5#3089A 251.21 0.0000 89.34 0.0000 71.69 0.0000 1.000240 ObjArcOff=0.001661,0.000000,4794723356.37000,0.001047,0.000000,4794723356.37000 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 02:22Z 11#3526 0.99998 198 1825 499 -14 56.5 6.1 15 226 --- 1.3" 05:40Z 6#3706B 1.00011 -1086 640 398 -61 65.0 5.6 16 261 --- 1.4" 06:54Z 8#3896A 1.00037 -3654 -1333 448 74 33.4 5.6 15 276 --- 10:46Z 9#3635B 1.00030 -2963 -825 505 130 76.6 4.1 23 245 --- 2.1" 11:35Z 5#3089A 1.00024 -2370 -393 455 251 89.4 4.1 19 264 --- 1.5" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:48Z 46 29 12 259 (W) 284 - 02:22Z 43 30 15 225 (SW) 497 - 05:41Z 42 30 16 261 (W) 418 - 06:54Z 42 30 15 276 (W) 368 - 10:47Z 39 28 22 245 (WSW) 296 - 11:35Z 39 29 19 264 (W) 383 - ---=== TELESCOPE STATUS ===--- 23:00Z - Began ventilation. Doors and louvers opened, fans on. 00:30Z - Opened. 12:05Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 12:28Z Telescope stowed at: 121,30,00 mount Instrument mounted: Cartridge 5 Counterweights at: 293 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro #2 141 lbs, 19 psi Spectro #1 135 lbs, 24 psi ---=== SOFTWARE USED ===--- ----STUI software STUI: 1.0b5 2010-10-18 alerts: trunk+svn114742 apo: v1_5 boss: v2_3_9 boss.daq: 1.2.14 boss.specMech: 3.0.1 gcamera: v1_2_3 guider: trunk+svn116784M hub: Tag: sdss3_v1_1_4 mcp: v6_6_5 mcp.plc: 25 platedb: v1_0_5 sop: v1_1_4 sos: v1_1_4 tcc: 2.13.0 -----Other software sdssProcedure(sdss3): v3_06 idlspec2d (sos3) --- v5_4_19 plate-mapper3 ------ NOCVS:v5_0_7 ------ IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_5 TPM: tpm_v3_5_0 AstroDa: v15_13_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.5210 -5.5100 Axial B 16.0920 16.0940 Axial C 15.9140 15.9220 Trans D 11.0010 11.0100 Lateral E 10.6680 10.6553 Lateral F 12.2555 12.2301 GALILS Commanded: 10600. -250. 5800. 10250. 13050. 12650. Actual: 10600. -250. 5800. 10250. 13053. 12647. SETMIR VALUES PriDesOrient: 0.00 -13.67 10.53 574.80 742.37 PriOrient: 0.00 -13.67 10.53 574.80 742.37 SECONDARY: ---------- Focus: 0.00 Air Temp.: 8.2 Alt.: 30.000467 MIGS TONIGHT NOMINAL Axial A -7.8350 -7.8320 Axial B -0.0430 -0.0410 Axial C 13.8160 13.8180 Trans D 14.1860 14.1840 GALILS Commanded: 1585325. 1566755. 1598880. -1450. -50. Actual: 1585350. 1566695. 1598880. -1462. -69. SETMIR VALUES SecDesOrient: 1257.00 -9.00 -4.00 -58.00 23.67 SecOrient: 1257.01 -8.99 -3.98 -58.07 23.31 ---=== PROBLEMS IN DETAIL ===--- Suggested faster procedure. 1. Perform collimation, calibration, guiding and field acquisition as usual 2. Take 2 science exposures with integration time of two minutes, guider on         # This will tell you how repeatable things are. 3. Turn off Guiding, continue tracking 4. Offset telescope by 0.5 arcseconds in RA from best guide position, take 1 science exposure with integration time of two minutes 5   clear offset and recenter with guider during readout 7. Offset telescope by -0.5 arcseconds in RA from best guide position, take 1 science exposure with integration time of two minutes 8   clear offset and recenter with guider during readout 9. Offset telescope by 0.5 arcseconds in DEC from best guide position, take 1 science exposure with integration time of two minutes 10    clear offset and recenter with guider during readout 11 Offset telescope by -0.5 arcseconds in DEC from best guide position, take 1 science exposure with integration time of two minutes     clear offset and recenter with guider during readout 12. Wait till guiding converges, take 2 science exposure with integration time of two minutes    # This will tell you how repeatable things are. 12. Turn off "automated focus" adjustment on guider, keep guiding 13. Move secondary mirror by +200 microns, take 1 science exposure with integration time of two minutes 14. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 15. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 16. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 17. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes - should be "best focus" position 18. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 19. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 20. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 21. Move secondary mirror by -50 microns, take 1 science exposure with integration time of two minutes 22  reset focus back to best focus, turn guider focus on 22. Wait till guiding/focus  converges, take 2 science exposure with integration time of two minutes    # This will tell you how repeatable things are. One question is what to do if the telescope scale changes during this, should put in between the 2 exposures at (12) and at (22)